OpenStudy (j.f):

Jenny's average score for six tests is 89. The teacher said that the lowest score would not be included in the average. If her lowest score was 79, what was the average of the remaining scores? brain fart:/

OpenStudy (mathstudent55):

How do you find a average test score? You add all test scores and divide by the number of tests. Ok so far?

OpenStudy (j.f):

yes

OpenStudy (mathstudent55):

In this case, you are told there were 6 tests and the average is 89. What is the sum of the test scores?

OpenStudy (mathstudent55):

\(average~score = \dfrac{sum~of~test~scores}{number~of~tests} \) \(89 = \dfrac{sum~of~test~scores}{6} \) In other words, you need to solve the equation above for "sum of test scores." To make it look simpler, let S = sum of test scores. Solve the simple equation below for S: \(89 = \dfrac{S}{6} \)

OpenStudy (mathstudent55):

Can you find S, the sum of the test scores?

OpenStudy (j.f):

534?

OpenStudy (mathstudent55):

Great. The sum of the scores of 6 tests is 534. Now we need the sum of the scores of the top 5 tests. The teacher said she can drop the lowest grade of 79. What will the sum of the top 5 scores be if you subtract 79 from 534?

OpenStudy (mathstudent55):

Then finally, since you will have the sum of the top 5 scores, all you need to do is divide that number by 5 to find the average of the top 5 scores.

OpenStudy (mathstudent55):

Sorry, gtg. Did you get an answer?

OpenStudy (j.f):

534-79=455 455/5=91

OpenStudy (mathstudent55):

Correct. Great job!

OpenStudy (anonymous):

correct...way to go. nice explaining mathstudent

OpenStudy (mathstudent55):

@kelliegirl33 Thanks.